# Conformal Distributed Remote Inference in Sensor Networks Under Reliability and Communication Constraints

### v1

@@ I-A Distributed Remote Inference
We focus on a scenario in which multiple sensors observe different features of the underlying input. For example, vision sensors may have different partial views of the scene of interest, and they may be subject to various random impairments, such as darkness, noise artifacts, and blurring. The goal is to perform multi-label classification tasks, such as binary segmentation, through communication to a central server.
@@ I-B Distributed Conformal Risk Control
To ensure deterministic, assumption-free reliability, distributed conformal prediction (D-CP) was introduced in [21] for settings without any communication constraints between sensors and central server. Focusing on conventional single-label classification problems, D-CP leverages online conformal prediction (CP) [22, 23] to dynamically adjust a shared local decision threshold used by all the sensors to construct a prediction set. D-CP ensures the deterministic coverage requirement that the true label is included in the prediction set with a target long-term rate. Furthermore, D-CP employs an online exponentiated gradient strategy [24] to sequentially adjust weights used by the central server to combine the sensors’ decisions, aiming to minimize the size of the prediction set.
@@ I-C Related Work
Conformal Risk Control: CP is a statistical framework that provides valid prediction intervals or sets with guaranteed coverage probability [25, 26]. CP has since been extended to distributed settings, particularly in federated scenarios, where data remains private across multiple devices while reliable decision-making is performed at the central server [27, 28]. These approaches operate offline and rely on data exchangeability. Online CP was developed in [22, 23] to maintain long-term calibration without statistical assumptions on data generation, ensuring the desired coverage frequency over time, even under adversarial data distribution shifts.
While CP focuses on maintaining coverage, conformal risk control (CRC) can address more general risk metrics, allowing control of risks such as FNR at any specified level [29, 30]. Online versions of CRC were introduced in [31, 32].
@@ II-A Setting
As depicted in Fig. 2, we consider a sensor network consisting of $K$ sensors and a central server, whose goal is to carry out distributed multi-label prediction under reliability and communication constraints. At each discrete time $t$, each sensor $k$ collects an input $X^{t}_{k}$ corresponding to a different noisy view of an underlying input $X^{t}$, which is associated with a label vector $Y^{t}\in\left\{0,1\right\}^{L}$. The goal is to use limited communication from sensors to the central server to support reliable detection of vector $Y^{t}$ at the server.
This setting captures multi-label classification tasks, in which there are $L$ possible labels describing the scene, e.g., $L$ possible objects, and the target variable $Y^{t}$ indicates which labels are pertinent to $X^{t}$. As shown in Fig. 1 and Fig. 2, this formulation encompasses the prediction of a target location in an image, in which each image $X^{t}$ with $L$ pixels is associated with a binary mask $Y^{t}\in\left\{0,1\right\}^{L}$, specifying the position of a target of interest. In particular, an element $Y^{t}_{l}=1$ indicates that the pixel $l$ will be part of the target, whereas $Y^{t}_{l}=0$ identifies the pixel $l$ as background. Practical applications include the prediction of the position of vehicles in images obtained from distributed cameras, e.g., for beamforming design in wireless networks [33]. Other examples include the prediction of wireless traffic [34] or spectrum sensing [35]. Henceforth, we will refer to each entry $Y^{t}_{l}$ of the vector $Y^{t}$ as a label.
@@
The weight $\beta^{t}_{k}$ reflects the contribution of the sensor $k$ to the global decision (3), and thus it should be ideally designed depending on the quality of the observation $X^{k}_{k}$ at sensor $k$. Finally, based on the soft estimate (3), using a global threshold $\theta^{t}$, the server obtains hard decisions for each label as
@@ II-B1 False negative rate and false positive rate
In general, the FNR measures the proportion of labels erroneously classified as irrelevant, i.e., $V^{t}_{k}=0$, while they are, in fact, pertinent, i.e., $Y^{t}_{k}=1$. As mentioned, in the segmentation example of Fig. 2, the FNR is the fraction of missed pixels containing the object of interest. The FNR is calculated as
@@ II-B2 Communication constraint
Each $L\times 1$ binary prediction vector $U^{t}_{k}$ in (2) must be communicated from sensor $k$ to the server. If no compression strategy is implemented, this requires the transmission of $L$ bits from each sensor. However, when there exist communication constraints between sensors and the server, this approach is generally inefficient. In fact, without any compression, sensors that do not obtain any useful information would require the same bandwidth as sensors that have non-trivial local decisions to communicate. Given that our main design goal is ensuring a constraint on the FNR (5), a local decision vector $U^{t}_{k}$ is deemed to be uninformative if it contains only ones. In fact, as discussed, in this case any FNR level is guaranteed, but without any discriminative power. Accordingly, we focus on compression strategies that require zero communicated bits when the decision corresponds to an all-one vector $U^{t}_{k}$, i.e.,
@@
since a threshold $\lambda^{t}_{k}\geq 1$ yields an all-zero decision $U^{t}_{k}$.
The overall communication cost from the $K$ sensors to the server at time $t$ is equal to the sum $\sum_{k=1}^{K}B^{t}_{k}\in[0,K]$.
@@ II-B4 Online optimization
We wish to address problem (9) in an online fashion by leveraging feedback or past decisions. Specifically, at each time $t$, upon making a decision $V^{t}$, the server receives feedback in the form of the ground-truth vector label $Y^{t}$. For instance, in the case of the prediction problem in Fig. 1, the decision $V^{t}$ may correspond to a prediction of the target’s position, which can be resolved at a later time. This feedback is used to calculate the long-term FNR using (5) and the long-term FPR using (6). Based on this information, the optimization variables $\left\{\beta_{k}^{t}\right\}_{k=1}^{K}$, $\left\{\lambda_{k}^{t}\right\}_{k=1}^{K}$, and $\theta^{t}$ are updated with the goal of addressing problem (9).
@@ III-B Combining Weights Update
$\Delta^{t}=\sum_{\tau=1}^{t}\left(\sum_{k=1}^{K}\beta^{\tau}_{k}P^{\tau}_{k}+\frac{1}{\eta^{\tau}}\ln\left(\sum_{k=1}^{K}\beta_{k}^{\tau}\cdot\exp\left\{-\eta^{\tau}P^{\tau}_{k}\right\}\right)\right).$ (15)
@@ IV Communication-Constrained Distributed Conformal Risk Control
This section proposes communication-constrained distributed CRC (CD-CRC), a novel protocol that addresses problem (9) in the presence of communication constraint. Unlike D-CRC, CD-CRC updates the local thresholds $\left\{\lambda^{t}_{k}\right\}_{k=1}^{K}$ in (2) by accounting for both constraints (9b) and (9c). However, the limitations imposed by the communication constraint (9c) prevent the updates of the local thresholds from being able to control the reliability requirement (9b) by themselves. To obviate this problem, DC-CRC integrates the local threshold updates with the global threshold update.
@@ IV-A Local Threshold Update
where $\beta^{t}_{k}$ are the weights used for combining in (3). The rationale for this choice is that the weight $\beta^{t}_{k}$ should be larger for sensors that have higher-quality observations. Note that other functions of the weights $\beta_{k}^{t}$ could also be used in lieu of (23), as long as a higher capacity is assigned to sensors $k$ with a larger value $\beta_{k}^{t}$.
@@ V-A Long-term FNR Constraint
Accordingly, as $T\rightarrow\infty$, the long-term FNR constraint in (9b) is satisfied.
@@ VI-A Setting
We employ pre-trained fully convolutional networks [41] with a ResNet-50 backbone [42], downloaded from the Torchvision repository. The last layer is modified to a convolutional layer followed by a softmax layer to produce the object probability vector $S^{t}_{k}$. The model is fine-tuned on the $2000$ training data points, using cross-entropy as the loss function. We employ the SGD optimizer with a learning rate of $0.005$, momentum of $0.9$, and weight decay of $0.0001$, over $50$ epochs.
@@
  • Blur effect ($X^{t}_{4}$): Sensor $4$ observes a random $60\times 60$ region of $X^{t}$ with a Gaussian blur with a kernel size of $7$.
@@
$C^{t}_{k}=\frac{C}{K},$ (36)
which is referred to as U-CD-CRC.
Figure 5: Time-averaged FNR (left), time-averaged communication load (middle), and time-averaged FPR (right) for CD-CRC, U-CD-CRC, and D-CRC versus the FNR constraint $\alpha$ with capacity constraint $C=1.5$.
Figure 6: Time-averaged global FNR (left), time-averaged communication load (middle), and time-averaged FPR (right) for CD-CRC, U-CD-CRC, and D-CRC versus the capacity constraint $C$ with FNR constraint $\alpha=0.2$.
@@ VI-B Performance Evaluation
To gain further insights into the operation of CD-CRC and D-CRC, the top-left corner of Fig. 4 shows the evolution of the weights $\left\{\beta_{k}^{t}\right\}_{k=1}^{K}$ over time $t$. It is observed that all schemes quickly identify sensor $1$ as the best. This way, CD-CRC allocates all communication resources to sensor $1$ via the adaptive resource allocation strategy (23), achieving local performance similar to D-CRC, which does not account for capacity constraint. CD-CRC meets the capacity constraint by disabling transmissions from non-optimal sensors, while, as shown in the bottom row of Fig. 4, D-CRC applies a common local threshold across all sensors, exceeding the capacity constraint.
@@
The impact of the capacity constraint $C$ with a fixed FNR constraint $\alpha=0.2$ is studied in Fig. 6. While CD-CRC and U-CD-CRC satisfy the capacity constraint for all values of $C$, D-CRC fails to do so when the capacity constraint is stringent (i.e., $C<2.5$). As compared with U-CD-CRC, CD-CRC with an adaptive capacity allocation requires a much smaller communication load by allocating all resources to the optimal sensor (sensor $1$), while disabling transmission from other non-optimal sensors. Moreover, when $C$ is insufficient even for a single sensor’s transmission, CD-CRC sacrifices the FPR of the optimal sensor to meet the capacity constraint. Conversely, when $C$ is large enough to support transmissions from all sensors, U-CD-CRC achieves the same FPR as CD-CRC and D-CRC.
@@ VII Conclusions
The theoretical analysis in this work offers insights into FNR and FPR worst-case guarantees, while simulation results demonstrate that CD-CRC outperforms existing methods, particularly in communication-constrained environments. Overall, CD-CRC presents a promising solution for reliable decision-making in distributed sensor networks under communication constraints, paving the way for further research in this area.
@@
Future work could explore the application of CD-CRC to more complex tasks such as binary segmentation with multiple objects or time-varying target processes with memory. Additionally, accounting for wireless impairments and investigating privacy-preserving technique, such as differential privacy, would be valuable directions for further study.
-A Proof of Positivity for the Corrected Global Threshold in CD-CRC
We begin by establishing a lower bound for the corrected global threshold $\tilde{\theta}^{t}$ in (25) following the Appendix A.1 in [31]. Suppose, for contradiction, that $\tilde{\theta}^{t}<\delta-\mu(1-\alpha)$ at some time $t$, while for all $t^{\prime}<t$, we have $\tilde{\theta}^{t^{\prime}}\geq\delta-\mu(1-\alpha)$. According to the update rule (25), we have
$\tilde{\theta}^{t-1}$ $=\tilde{\theta}^{t}+\mu\left(N^{t-1}-\alpha\right)$
$\leq\tilde{\theta}^{t}+\mu\left(1-\alpha\right)<\delta.$ (37)
Given that $N^{t-1}=0\leq\alpha$ when $\tilde{\theta}^{t-1}<\delta$ ($\theta^{t-1}<0$), we have
$\tilde{\theta}^{t}$ $=\tilde{\theta}^{t-1}-\mu\left(N^{t-1}-\alpha\right)$
$\geq\tilde{\theta}^{t-1}\geq\delta-\mu(1-\alpha),$ (38)
which contradicts the initial assumption. Thus, $\tilde{\theta}^{t}$ must satisfy
$\tilde{\theta}^{t}\geq\delta-\mu(1-\alpha)>0,$ (39)
where the last inequality uses the setting (27).
-B Proof of Theorem 2
By substituting $t=1,2,\ldots,T$ into the update rule in (25) and by summing up the contributions over time $t$, the average global FNR over time $T$ can be upper bounded as
$\frac{1}{T}\sum_{t=1}^{T}N^{t}$ $=\alpha+\frac{\tilde{\theta}^{1}-\tilde{\theta}^{T+1}}{\mu T}$
$\leq\alpha+\frac{\tilde{\theta}^{1}-\delta+\mu\left(1-\alpha\right)}{\mu T}$
$<\alpha+\frac{\tilde{\theta}^{1}}{T},$ (40)
where the inequalities use (39).
-C Proof of Theorem 3
Following Appendix A.1 in [31], we start by providing an upper bound of the local threshold $\lambda^{t}_{k}$ in the following lemma.
Lemma 2.
For any time $t$ and sensor $k$, the local threshold satisfies the inequality $\lambda_{k}^{t}\geq-2\gamma$, where $\gamma$ denotes the step size in the update rule (IV-A).
Proof: Assume, contradicting the desired inequality, that there exists a time $t$ such that $\lambda_{k}^{t}<-2\gamma$. Further assume that for all $t^{\prime}<t$, we have $\lambda^{t^{\prime}}_{k}\geq-2\gamma$. According to (IV-A) and given that $B^{t-1}_{k}\in[0,1]$, $C^{t-1}_{k}\in[0,C]$, $N^{t-1}\in[0,1]$, and $\alpha\in[0,1]$, we have
$\lambda_{k}^{t-1}$ $=\lambda^{t}_{k}+\gamma\Delta\lambda^{t-1}_{k}$
$\leq\lambda^{t}_{k}+\gamma<-\gamma<0.$ (41)
Based on the fact that $B^{t-1}_{k}=0\leq C^{t-1}_{k}$ when $\lambda_{k}^{t-1}<0$, we further consider the following two situations.
If $\lambda_{k}^{t-1}\geq-\gamma$, we have
$\lambda^{t}_{k}$ $=\lambda_{k}^{t-1}-\gamma\max\left\{N^{t-1}-\alpha,B^{t-1}_{k}-C^{t-1}_{k}\right\}$
$\geq\lambda^{t-1}_{k}-\gamma\geq-2\gamma,$ (42)
which contradicts our initial assumption.
If $\lambda_{k}^{t-1}<-\gamma$, we have
$\lambda^{t}_{k}$ $=\lambda_{k}^{t-1}\geq-2\gamma,$ (43)
which again contradicts our initial assumption.
We now prove Theorem 3. According to (IV-A), we have the inequality
$\Delta\lambda^{t}_{k}\geq B^{t}_{k}-C^{t}_{k}.$ (44)
By substituting $t=1,2,\ldots,T$ and $k=1,2\ldots,K$ into the update rule in (IV-A) and by summing up all contributions over time $t$, the average overall required communication costs for all sensors over time $T$ can be expressed as
$\frac{1}{T}\sum_{t=1}^{T}\sum_{k=1}^{K}B^{t}_{k}$ $\leq C+\sum_{t=1}^{T}\sum_{k=1}^{K}\Delta\lambda^{t}_{k}$
$=C+\frac{\sum_{k=1}^{K}\left(\lambda^{1}_{k}-\lambda^{T+1}_{k}\right)}{\gamma T}$
$\leq C+\frac{\sum_{k=1}^{K}\left(\lambda^{1}_{k}+2\gamma\right)}{\gamma T},$ (45)
where the first inequality uses (44) and the third inequality uses Lemma 2.
-D Proof of Lemma 1
This proof generalizes the proof of Theorem 2.14 in [21]. By substituting (3) and (26) into (4), the $l$-th entry in $V^{t}$ can be expressed as
$V^{t}_{l}=\mathds{1}\left\{\sum_{k=1}^{K}\beta_{k}^{t}U_{k,l}^{t}+\delta\geq\tilde{\theta}^{t}\right\}.$ (46)
The global FPR in (6) can be upper bounded as
$P^{t}$ $=\frac{\sum_{l=1}^{L}\left(1-Y^{t}_{l}\right)\mathds{1}\left\{\sum_{k=1}^{K}\beta_{k}^{t}U_{k,l}^{t}+\delta\geq\tilde{\theta}^{t}\right\}}{\sum_{l=1}^{L}\left(1-Y^{t}_{l}\right)}$
$=\frac{\sum_{l=1}^{L}\left(1-Y^{t}_{l}\right)\mathds{1}\left\{1/\tilde{\theta}^{t}\sum_{k=1}^{K}\beta_{k}^{t}U_{k,l}^{t}+\delta/\tilde{\theta}^{t}\geq 1\right\}}{\sum_{l=1}^{L}\left(1-Y^{t}_{l}\right)}$
$\leq\frac{1}{\tilde{\theta}^{t}}\frac{\sum_{l=1}^{L}\left(1-Y^{t}\right)\sum_{k=1}^{K}\beta_{k}^{t}U_{k,l}^{t}}{\sum_{l=1}^{L}\left(1-Y^{t}_{l}\right)}+\frac{\delta}{\tilde{\theta}^{t}}$
$=\frac{1}{\tilde{\theta}^{t}}\sum_{k=1}^{K}\beta^{t}_{k}\frac{\sum_{l=1}^{L}\left(1-Y^{t}\right)U^{t}_{k,l}}{\sum_{l=1}^{L}\left(1-Y^{t}_{l}\right)}+\frac{\delta}{\tilde{\theta}^{t}}$
$=\frac{1}{\tilde{\theta}^{t}}\sum_{k=1}^{K}\beta^{t}_{k}P^{t}_{k}+\frac{\delta}{\tilde{\theta}^{t}},$ (47)
where the third inequality uses the fact that $\mathds{1}\left\{x\geq 1\right\}\leq x$ for any $x\geq 0$.
-E Proof of Theorem 4
We begin by determining the range of the corrected global threshold $\tilde{\theta}^{t}$ in CD-CRC. By applying a contradiction method similar to that used in Appendix -A, we can establish the following bounds,
$\delta-\mu(1-\alpha)\leq\tilde{\theta}^{t}\leq 1+\delta+\mu\alpha.$ (48)
As a result, the average global FPR can be upper bounded as
$\frac{1}{T}\sum_{t=1}^{T}P^{t}\leq\frac{1}{T}\sum_{t=1}^{T}\sum_{k=1}^{K}\beta^{t}_{k}\frac{P^{t}_{k}}{\tilde{\theta}^{t}}+\frac{1}{T}\sum_{t=1}^{T}\frac{\delta}{\tilde{\theta}^{t}}$
$\leq$ $\frac{2}{T}\left(\ln K\cdot\frac{\max\limits_{t\in\left\{1,\ldots,T\right\}}\Delta P^{t}/\tilde{\theta}^{t}}{\sum_{t=1}^{T}\Delta P^{t}/\tilde{\theta}^{t}}\right)^{\frac{1}{2}}\cdot\left(\sum_{t=1}^{T}\frac{P^{t}_{\textrm{max}}}{\tilde{\theta}^{t}}-\min_{k\in\left\{1,\ldots,K\right\}}\sum_{t=1}^{T}\frac{P_{k}^{t}}{\tilde{\theta}^{t}}\right)^{\frac{1}{2}}$
$\cdot\left(\min_{k\in\left\{1,\ldots K\right\}}\sum_{t=1}^{T}\frac{P_{k}^{t}}{\tilde{\theta}^{t}}-\sum_{t=1}^{T}\frac{P^{t}_{\textrm{min}}}{\tilde{\theta}^{t}}\right)^{\frac{1}{2}}$
$+\frac{1}{T}\left(\frac{16}{3}\ln K+2\right)\max_{t\in\left\{1,\ldots,T\right\}}\frac{\Delta P^{t}}{\tilde{\theta}^{t}}+\frac{1}{T}\min_{k}\sum_{t=1}^{T}\frac{P_{k}^{t}}{\tilde{\theta}^{t}}+\frac{1}{T}\sum_{t=1}^{T}\frac{\delta}{\tilde{\theta}^{t}}$
$\leq$ $\frac{2}{T}\left(\frac{\left(1+\delta+\mu\alpha\right)\ln K}{\delta-\mu(1-\alpha)}\cdot\frac{\max\limits_{t\in\left\{1,\ldots,T\right\}}\Delta P^{t}}{\sum_{t=1}^{T}\Delta P^{t}}\right)^{\frac{1}{2}}$
$+\frac{1}{T\left(\delta-\mu(1-\alpha)\right)}\left(\frac{16}{3}\ln K+2\right)\max\limits_{t\in\left\{1,\ldots,T\right\}}\Delta P^{t}$
$+\frac{\mathcal{P}^{*}}{\delta-\mu(1-\alpha)}+\frac{1}{T}\sum_{t=1}^{T}\frac{\delta}{\tilde{\theta}^{t}},$ (49)
where the first inequality uses Lemma 1, the second inequality extends from Theorem 1 (see also Theorem 8 in [24]), and the third inequality uses (48).

### v3

@@ I-A Distributed Remote Inference
We focus on a scenario in which multiple sensors observe different features of the underlying input. For example, vision sensors may have different partial views of the scene of interest, and they may be subject to various random impairments, such as darkness, noise artifacts, and blurring. The goal is to perform multi-label classification tasks–where each instance can be associated with multiple labels simultaneously, such as binary segmentation–through communication to a central server.
@@ I-B Distributed Conformal Risk Control
To ensure deterministic, assumption-free reliability, distributed conformal prediction (D-CP) was introduced in [21] for settings without any communication constraints between sensors and central server. Focusing on conventional single-label classification problems, D-CP leverages online conformal prediction (CP) [22] to dynamically adjust a shared local decision threshold used by all the sensors to construct a prediction set. D-CP ensures the deterministic coverage requirement that the true label is included in the prediction set with a target long-term rate. Furthermore, D-CP employs an online exponentiated gradient strategy [23] to sequentially adjust weights used by the central server to combine the sensors’ decisions, aiming to minimize the size of the prediction set.
@@ I-C Related Work
Conformal Risk Control: CP is a statistical framework that provides valid measures of predictive uncertainty by generating prediction intervals or sets that contain the true label with a specific probability [24, 25]. CP has since been extended to distributed settings, particularly in federated scenarios, where data remains private across multiple devices while reliable decision-making is performed at the central server [26, 27]. These approaches operate offline and rely on data exchangeability. Online CP, developed in [22] and further studied in [28, 29, 30], maintains long-term calibration without statistical assumptions on data generation. This way, online CP ensures the desired coverage frequency over time, even under adversarial data distribution shifts.
While CP focuses on maintaining coverage, conformal risk control (CRC) can address more general risk metrics, allowing control of risks such as FNR at any specified level [31, 32, 33]. Online versions of CRC were introduced in [34, 35].
@@ II-A Setting
As depicted in Fig. 2, we consider a sensor network consisting of $K$ sensors and a central server, whose goal is to carry out distributed multi-label prediction under reliability and communication constraints. Specifically, at each discrete time $t$, the system operates as follows:
  • 1.
    The sensors send information to the central server.
  • 2.
    The central server addresses a multi-label classification problem, such as segmentation. Depending on the application, the server may act on the classification decision. For instance, in a wireless communication setting, the central server may use the segmentation decision to direct the antenna beams towards the mobile users [36].
  • 3.
    The central server obtains feedback on the false negative rate and false positive rate performance of its classification decision, which will be formally defined below. This feedback may be obtained as a result of the actions taken based on the classification decisions. For example, in the mentioned wireless use case, the server may be informed about the fraction of users that did not receive timely service.
In more detail, at each discrete time $t$, each sensor $k$ collects an input $X^{t}_{k}$ corresponding to a different noisy view of an underlying input $X^{t}$, which is associated with a label vector $Y^{t}\in\left\{0,1\right\}^{L}$. The goal is to use limited communication from sensors to the central server to support reliable detection of vector $Y^{t}$ at the server.
This setting captures multi-label classification tasks, in which there are $L$ possible labels describing the scene, e.g., $L$ possible objects, and the target variable $Y^{t}$ indicates which labels are pertinent to $X^{t}$. As shown in Fig. 1 and Fig. 2, this formulation encompasses the segmentation of an image, in which each image $X^{t}$ with $L$ pixels is associated with a binary mask $Y^{t}\in\left\{0,1\right\}^{L}$ specifying the position of an object of interest. In particular, an element $Y^{t}_{l}=1$ indicates that the pixel $l$ is part of the target, whereas $Y^{t}_{l}=0$ identifies the pixel $l$ as background. Henceforth, we will refer to each entry $Y^{t}_{l}$ of the vector $Y^{t}$ as a label.
@@
The weight $\beta^{t}_{k}$ reflects the contribution of the sensor $k$ to the global decision (3), and thus it should be ideally designed depending on the quality of the observation $X^{t}_{k}$ at sensor $k$. Finally, based on the soft estimate (3), using a global threshold $\theta^{t}$, the server obtains hard decisions for each label as
@@ II-B1 False negative rate and false positive rate
In general, the FNR measures the proportion of labels erroneously classified as irrelevant, i.e., $V^{t}_{l}=0$, while they are, in fact, pertinent, i.e., $Y^{t}_{l}=1$ for label $l$. As mentioned, in the segmentation example of Fig. 2, the FNR is the fraction of missed pixels containing the object of interest. The FNR is calculated as
@@ II-B2 Communication constraint
Each $L\times 1$ binary prediction vector $U^{t}_{k}$ in (2) must be communicated from sensor $k$ to the server. If no compression strategy is implemented, this requires the transmission of $L$ bits from each sensor. However, when there exist communication constraints between sensors and the server, this approach is generally inefficient. In fact, without any compression, sensors that do not obtain any useful information would require the same bandwidth as sensors that have non-trivial local decisions to communicate. Given that our main design goal is ensuring a constraint on the FNR (5), a local decision vector $U^{t}_{k}$ is deemed to be uninformative if it contains only ones. In fact, as discussed, in this case any FNR level is guaranteed, but without any discriminative power.
Accordingly, we focus on compression strategies that require zero communicated bits when the decision corresponds to an all-one vector $U^{t}_{k}$, i.e.,
@@
since a threshold $\lambda^{t}_{k}\geq 1$ yields an all-zero decision $U^{t}_{k}$. Accordingly, the communication cost per sensor refers to the average number of bits per element of the decision vector. Moreover, the overall communication cost from the $K$ sensors to the server at time $t$ is equal to the sum $\sum_{k=1}^{K}B^{t}_{k}\in[0,K]$.
@@ II-B4 Online optimization
We wish to address problem (9) in an online fashion by leveraging feedback or past decisions. Specifically, at each time $t$, upon making a decision $V^{t}$, the server receives feedback about the FNR (5) and the FPR (6). Importantly, feedback is collected only after a decision is made, and only about the values of the FNR and the FPR, not about the original multi-label $Y^{t}$. Based on the received feedback, the optimization variables $\left\{\beta_{k}^{t}\right\}_{k=1}^{K}$, $\left\{\lambda_{k}^{t}\right\}_{k=1}^{K}$, and $\theta^{t}$ are updated with the goal of addressing problem (9).
The assumed feedback framework captures many settings of practical interest. For example, in wireless communication, segmentation decisions used to adjust antenna beams toward mobile users can be assessed based on feedback from the users [36]. As another application, in security and surveillance use cases, threat detection decisions can be evaluated in hindsight using feedback from human inspection or forensic evidence [37].
As a final remark, as we will further discuss in Sec. V, meeting the FNR requirement (9b) does not require the exact feedback of the FNR. In fact, a conservative estimate of the FNR, i.e., an upper bound on the FNR (5), is sufficient to ensure the reliability requirement (9b).
@@ III-B Combining Weights Update
$\Delta^{t}=\sum_{\tau=1}^{t}\left(\sum_{k=1}^{K}\beta^{\tau}_{k}P^{\tau}_{k}-\left(-\frac{1}{\eta^{\tau}}\ln\left(\sum_{k=1}^{K}\beta_{k}^{\tau}\cdot\exp\left\{-\eta^{\tau}P^{\tau}_{k}\right\}\right)\right)\right).$ (15)
@@ IV Communication-Constrained Distributed Conformal Risk Control
This section proposes communication-constrained distributed CRC (CD-CRC), a novel protocol that addresses problem (9) in the presence of communication constraint. By design, CD-CRC inherits the performance guarantees of D-CRC when the channel capacity is sufficiently large to make the capacity constraint (9c) immaterial, while being able to address problem (9) even in the presence of stringent capacity constraints.
Unlike D-CRC, CD-CRC updates the local thresholds $\left\{\lambda^{t}_{k}\right\}_{k=1}^{K}$ in (2) by accounting for both constraints (9b) and (9c). However, the limitations imposed by the communication constraint (9c) prevent the updates of the local thresholds from being able to control the reliability requirement (9b) by themselves. To obviate this problem, DC-CRC integrates the local threshold updates with the global threshold update.
@@ IV-A Local Threshold Update
where $\beta^{t}_{k}$ are the weights used for combining in (3). The rationale for this choice is that the weight $\beta^{t}_{k}$ should be larger for sensors that have higher-quality observations. Note that other functions of the weights $\beta_{k}^{t}$ could also be used in lieu of (23), as long as a higher capacity is assigned to sensor $k$ with a larger value $\beta_{k}^{t}$.
@@ V-A Long-term FNR Constraint
According to (31), as $T\rightarrow\infty$, the long-term FNR constraint in (9b) is satisfied. As anticipated in Sec. II-B4, the result (31) can be extended to a setting with imperfect FNR feedback. To elaborate, suppose that, at each time $t$, the central server is only given a conservative estimate $\hat{N}^{t}\geq N^{t}$ of the FNR in (5). Replacing $\hat{N}^{t}$ for $N^{t}$ in the update (25), CD-CRC also guarantees the condition (31), as it follows directly from the proof in Appendix -B.
@@ VI-A Setting
We employ pre-trained fully convolutional networks [43] with a ResNet-50 backbone [44], downloaded from the Torchvision repository. The last layer is modified to a convolutional layer followed by a softmax layer to produce the object probability vector $S^{t}_{k}$. The model is fine-tuned on the $2000$ training data points using cross-entropy as the loss function. We employ the SGD optimizer with a learning rate of $0.005$, momentum of $0.9$, and weight decay of $0.0001$ over $50$ epochs.
@@
  • Blur effect ($X^{t}_{4}$): Sensor $4$ observes a random $60\times 60$ region of $X^{t}$ with a Gaussian blur of kernel size $7$.
For additional experiments conducted under different settings, please refer to Appendix -F.
@@
$C^{t}_{k}=\frac{C}{K}.$ (36)
The latter, referred to as U-CD-CRC, serves as a natural benchmark to validate the effectiveness of the proposed capacity allocation strategy (23) for CD-CRC.
Figure 5: Time-averaged FNR (left), time-averaged communication load (middle), and time-averaged FPR (right) for CD-CRC, U-CD-CRC, and D-CRC [21] versus the FNR constraint $\alpha$ with capacity constraint $C=1.5$.
Figure 6: Time-averaged FNR (left), time-averaged communication load (middle), and time-averaged FPR (right) for CD-CRC, U-CD-CRC, and D-CRC [21] versus the capacity constraint $C$ with FNR constraint $\alpha=0.2$.
@@ VI-B Performance Evaluation
To gain further insights into the operation of CD-CRC and D-CRC, the top-left corner of Fig. 4 shows the evolution of the weights $\left\{\beta_{k}^{t}\right\}_{k=1}^{K}$ over time $t$. It is observed that all schemes quickly identify sensor $1$ as the best due to the online exponentiated gradient rules in (12) and (28), which rapidly assign full weight to the sensor due to its superior FPR performance. This way, CD-CRC allocates all communication resources to sensor $1$ via the adaptive resource allocation strategy (23) which allocates communication resources proportionally to the sensors’ weights, achieving local performance similar to D-CRC, which does not account for capacity constraint. CD-CRC meets the capacity constraint by disabling transmissions from non-optimal sensors, while, as shown in the bottom row of Fig. 4, D-CRC applies a common local threshold across all sensors, exceeding the capacity constraint.
@@
Based on the analysis in Fig. 4 and Fig. 5, CD-CRC can achieve the same ideal performance of D-CRC, while also satisfying the capacity constraints by efficiently utilizing the available communication resources. In fact, unlike U-CD-CRC, which evenly distributes capacity, and D-CRC, which ignores capacity constraints, CD-CRC adaptively allocates capacity across the sensors, avoiding unnecessary transmissions and adhering to the capacity constraint.
The impact of the capacity constraint $C$ with a fixed FNR constraint $\alpha=0.2$ is studied in Fig. 6. While CD-CRC and U-CD-CRC satisfy the capacity constraint for all values of $C$, D-CRC fails to do so when the capacity constraint is stringent (i.e., $C<2.5$). As compared with U-CD-CRC, CD-CRC with an adaptive capacity allocation requires a much smaller communication load by allocating all resources to the optimal sensor (sensor $1$), while disabling transmission from other non-optimal sensors. Moreover, when $C$ is insufficient even for a single sensor’s transmission (i.e., $C<1$), CD-CRC sacrifices the FPR of the optimal sensor to meet the capacity constraint. Conversely, when $C$ is large enough to support transmissions from all sensors, U-CD-CRC achieves the same FPR as CD-CRC and D-CRC.
@@ VII Conclusions
The theoretical analysis in this work offers insights into FNR and FPR worst-case guarantees, while simulation results demonstrate that CD-CRC outperforms existing methods, particularly in communication-constrained environments. Overall, CD-CRC presents a promising solution for reliable decision-making in distributed sensor networks under communication constraints, paving the way for further research in this area. Additionally, the CD-CRC framework can be seamlessly integrated with online model updates [39, 46], allowing for enhanced adaptability and performance in applications requiring real-time model adjustments. We leave a full investigation of this aspect to future work.
@@
Future work could explore the application of CD-CRC to more complex tasks such as binary segmentation with multiple objects or time-varying target processes with memory. Additionally, accounting for wireless impairments and investigating privacy-preserving techniques, such as differential privacy, would be valuable directions for further study.